OpenStudy (anonymous):

Evaluate: [(30 + 6) − 32] ÷ 9 • 2

OpenStudy (anonymous):

PEMDAS

OpenStudy (anonymous):

i keep doing some thing wrong @hereweg0

OpenStudy (anonymous):

ok the (36-32)/9*2 (4/9)*2 8/9

OpenStudy (anonymous):

36-32= 4/18 = 1 /4

OpenStudy (anonymous):

1) Follow PEMDAS so parentheses/brackets come first 2) So you do 30+6 to get 36 and it is in a bracket so your not yet 3) Now the expression looks like [36-32]/9*2 4) You finish whatever is in the brackets which makes the expression 4/9*2 5) The next part is to multiply or divide and you have to do whatever somes first, which is division 6) 4/9 can stay like that and just multiply it with 2 to get 8/9 as a final answer hope that helps :)

OpenStudy (anonymous):

But when you do PEMDAS you have to multiply 9*2

OpenStudy (anonymous):

no its interchangeable, it is which ever one comes first, you can evalute this expression on a scientific calculator and get my answer

OpenStudy (anonymous):

Ooo Kay